# On the Real Zeroes of Half-integral Weight Hecke Cusp Forms, II

### v1

@@ 6.1. Random model
We need a probabilistic random model for the Hecke eigenvalues $\lambda_{f}(m)$, which we now introduce following [19]. To motivate the model, let $G^{\sharp}$ denote the set of conjugacy classes of $G=\mathrm{SU}(2)$ endowed with the Sato–Tate measure. Note that if $m>1$ has the prime factorisation $m=p_{1}^{\alpha_{j}}\cdots p_{\ell}^{\alpha_{\ell}}$, then we have
@@
with $\alpha_{p},\beta_{p}$ being the Satake parameters of $f$ at a prime $p$. Then it can be shown that for a fixed set of primes $\{p_{1},...,p_{\ell}\}$, the $\ell$-tuple $\{g_{f}^{\sharp}(p_{1}),...,g_{f}^{\sharp}(p_{\ell})\}_{f\in H_{2k}}$ of conjugacy classes equidistributes inside $(G^{\sharp})^{\ell}$ as $k\longrightarrow\infty$.
This suggest the following random model for $\lambda_{f}(m)$. Let $\{g_{p}^{\sharp}\}_{p}$ be a sequence of independent random variables with values in $G^{\sharp}$ that are distributed according to the Sato–Tate measure. Then we define $X(1)=1$ and
@@
and $\mathbb{E}(X(m))=1_{m=0}$. By linearity of the expectation these give
(6.2) $\mathbb{E}(X(m)X(n))=\sum_{\begin{subarray}{c}d|(m,n)\\ mn=d^{2}\end{subarray}}1=\begin{cases}1&\text{if }m=n\\ 0&\text{otherwise}\end{cases}$
@@ 6.2. Construction of a random mollifier
$M_{1}(X;d)=(\log K)^{1/2}\sum_{n\leq K^{2\theta_{J}}}\frac{\lambda(n)\chi_{d}(n)h_{1}(n)}{2^{\Omega(n)}\sqrt{n}}\sum_{u|n}c_{n}(u)X(u).$
@@ 7. Character sum
Let $d$ be an odd squarefree integer and let $c,v,$ and $\eta$ be natural numbers so that $v\eta=u[c,d]^{2}/c^{2}$. Then the sum
$\sum_{x\,([c,d])}\sum_{w\,([c,d])}\chi_{d}(x)\chi_{d}(w)S\left(xw,v\eta;c\right)e\left(\frac{xv+w\eta}{[c,d]}\right)$
@@
(7.1) $\mathop{{\sum}^{\makebox[0.0pt][l]{$\star$}}}_{\gamma\,(c)}\sum_{w\,([c,d])}\chi_{d}(w)e\left(\frac{v\eta\overline{\gamma}}{c}+\frac{w\eta}{[c,d]}\right)\sum_{x\,([c,d])}\chi_{d}(x)e\left(\frac{xw\gamma}{c}+\frac{xv}{[c,d]}\right).$
@@
Suppose that $n$ and $u$ are odd positive integers with $nu\leq X^{1-\varepsilon}$. Let $\Phi$ denote a smooth and compactly supported function in $[1/2,5/2]$. Then
$\mathop{{\sum}^{\raisebox{-3.0pt}{\makebox[0.0pt][l]{$\flat$}}}}_{d}\chi_{d}(nu)\Phi\left(\frac{|d|}{X}\right)V_{k}\left(\frac{u}{|d|}\right)=\frac{X\pi^{2}}{9}\left(\int_{0}^{\infty}\Phi(\xi)\,\mathrm{d}\xi\right)\prod_{p|nu}\left(1+\frac{1}{p}\right)^{-1}\cdot 1_{nu=\square}+O_{\varepsilon}\left(X^{3/4}k^{-1/4+\varepsilon}+X^{1/2+\varepsilon}\sqrt{nu}\right).$
@@
We will evaluate the $r$-sum by applying Lemma 5.7. The terms where $nu$ is a square will contribute the main term in the zero frequency term on the dual side and the rest will give the error term.
@@
(8.3) $\frac{X}{2}\cdot\frac{1}{2\pi i}\int_{(c)}\left(\int_{0}^{\infty}\Phi(\xi)\xi^{s}\,\mathrm{d}\xi\right)\frac{\Gamma(s+k)}{\Gamma(k)}\left(\frac{X}{2\pi}\right)^{s}e^{s^{2}}\sum_{\begin{subarray}{c}\alpha=1\\ (\alpha,2u)=1\end{subarray}}^{\infty}\frac{\mu(\alpha)}{\alpha^{2}}u^{-s}\frac{\varphi(nu)}{nu}\cdot 1_{nu=\square}\,\frac{\mathrm{d}s}{s}.$
@@
(8.5) $\frac{X}{4}\prod_{p\neq 2}\left(1-\frac{1}{p^{2}}\right)\prod_{p|nu}\left(1+\frac{1}{p}\right)^{-1}\cdot 1_{nu=\square}\cdot\frac{1}{2\pi i}\int_{(c)}\left(\int_{0}^{\infty}\Phi(\xi)\xi^{s}\,\mathrm{d}\xi\right)\frac{\Gamma(s+k)}{\Gamma(k)}\left(\frac{X}{2\pi}\right)^{s}e^{s^{2}}u^{-s}\,\frac{\mathrm{d}s}{s}.$
@@
$\frac{X\pi^{2}}{9}\left(\int_{0}^{\infty}\Phi(\xi)\,\mathrm{d}\xi\right)\prod_{p|nu}\left(1+\frac{1}{p}\right)^{-1}\cdot 1_{nu=\square}+O_{\varepsilon}\left(X^{3/4}k^{-1/4+\varepsilon}\right),$
@@
It remains to bound the contribution coming from the other terms (i.e. $j\neq 0$) after applying Lemma 5.7. Let us set $Y:=X^{1/2-\varepsilon}/\sqrt{nu}$ and
@@
We split the $\alpha$-sum in (8.2) into two parts corresponding to $\alpha\leq Y$ and $\alpha>Y$. For the latter terms we estimate the $r$-sum trivially and get the upper bound (recall that $nu\leq X^{1-\varepsilon}$ and $r\alpha^{2}\asymp X$ so $V_{k}(u/r\alpha^{2})\ll 1$ uniformly in $k$ in this range)
For the terms with $\alpha\leq Y$ the sum we have to estimate after an application of Poisson summation is
@@
$=(\log K)^{1/2}\sum_{k\in\mathbb{Z}}h\left(\frac{2k-1}{K}\right)\sum_{f\in B_{k}}\omega_{f}\mathop{{\sum}^{\raisebox{-3.0pt}{\makebox[0.0pt][l]{$\flat$}}}}_{d}\,L\left(\frac{1}{2},f\otimes\chi_{d}\right)\phi\left(\frac{|d|}{X}\right)\sum_{n\leq K^{2\theta_{J}}}\frac{h_{1}(n)a_{f}(n)\lambda(n)\chi_{d}(n)}{2^{\Omega(n)}\sqrt{n}}.$
Let us first use the approximate functional equation and then execute the $f$-sum. Using the definition of $a_{f}(n)$ and applying the Petersson formula we have
@@
where the error term
@@
$\qquad\qquad\times\sum_{n\leq K^{2\theta_{J}}}\frac{h_{1}(n)\lambda(n)\chi_{d}(n)}{2^{\Omega(n)}\sqrt{n}}\sum_{u|n}c_{n}(u)\sum_{c=1}^{\infty}\frac{S(m,u;c)}{c}J_{2k-1}\left(\frac{4\pi\sqrt{mu}}{c}\right)$
gives a negligible contribution estimating trivially using Weil’s bound for Kloosterman sums and the bound (5.16) as $mu\leq mn\ll K^{1+2\theta_{J}+\varepsilon}X$ by recalling that $J$ was chosen so that $\eta_{2}\leq\theta_{J}\leq e\eta_{2}$ for some very small $\eta_{2}>0$ as well as $X\ll\sqrt{K}$.
Thus the main term is given by
$(\log K)^{1/2}\sum_{k\in\mathbb{Z}}h\left(\frac{2k-1}{K}\right)\mathop{{\sum}^{\raisebox{-3.0pt}{\makebox[0.0pt][l]{$\flat$}}}}_{d}\phi\left(\frac{|d|}{X}\right)\sum_{n\leq K^{2\theta_{J}}}\frac{h_{1}(n)\lambda(n)\chi_{d}(n)}{2^{\Omega(n)}\sqrt{n}}\sum_{u|n}c_{n}(u)\frac{\chi_{d}(u)}{\sqrt{u}}V_{k}\left(\frac{u}{|d|}\right).$
@@
after executing the $k$-sum using Poisson summation as the error term from Lemma 8.1 contributes $\ll_{\varepsilon}K^{3/4+\varepsilon}X^{3/4}+K^{1-\varepsilon}X^{1/2}$.
@@
$\qquad\qquad=\prod_{p\in I_{j}}\left(1-\frac{\iota(p)\nu_{2}(p)}{2\sqrt{p}}\left(1+\frac{1}{p}\right)^{-1}+\frac{\iota(p^{2})\nu_{2}(p^{2})}{2p}\left(1+\frac{1}{p}\right)^{-1}+O\left(\frac{1}{p^{3/2}}\right)\right).$
Combining the previous estimates with the fact that
$\prod_{p\in I_{j}}\left(1-\frac{\iota(p)\nu_{2}(p)}{2\sqrt{p}}\left(1+\frac{1}{p}\right)^{-1}+\frac{\iota(p^{2})\nu_{2}(p^{2})}{2p}\left(1+\frac{1}{p}\right)^{-1}+O\left(\frac{1}{p^{2}}\right)\right)^{-1}\ll 1_{j=0}\cdot(\log K)^{O(1)}+1$
@@
$XK(\log K)^{1/2}\prod_{c<p\leq K^{\theta_{J}}}\left(1-\frac{\iota(p)\nu_{1}(p)}{2\sqrt{p}}\left(1+\frac{1}{p}\right)^{-1}+\frac{\iota(p^{2})\nu_{1}(p^{2})}{2p}\left(1+\frac{1}{p}\right)^{-1}+O\left(\frac{1}{p^{2}}\right)\right)$
@@
To estimate the second product from below note that
@@
as $\eta_{1}$ was chosen to be sufficiently small.
@@
$\prod_{c<p<K^{\theta_{J}}}\left(1-\frac{1}{2p}+O\left(\frac{1}{p^{2}}\right)\right)\gg(\log K)^{-1/2}.$
@@ 9. Relating the Fourth moment to a random model
Let $n\leq K^{\theta_{J}}$ be a positive integer. Then we have
@@
$=\mathop{{\sum}^{\raisebox{-3.0pt}{\makebox[0.0pt][l]{$\flat$}}}}_{d}\phi\left(\frac{|d|}{X}\right)\chi_{d}(n)\sum_{k\sim K}\mathbb{E}\left(L(X;d,k)\sum_{u|n}c_{n}(u)X(u)\right)$
$\qquad+2K\widehat{h}(0)\mathop{{\sum}^{\raisebox{-3.0pt}{\makebox[0.0pt][l]{$\flat$}}}}_{d}\,\frac{\varphi(d)}{d}\phi\left(\frac{d}{X}\right)\sum_{u|n}\frac{c_{n}(u)}{\sqrt{u}}\chi_{d}(nu)\cdot\frac{1}{2\pi i}\int_{(\varepsilon)}\frac{e^{2x^{2}}}{x^{2}}\cdot u^{-x}\sum_{v|u}v^{2x}\,\mathrm{d}x+O\left(X^{3/4}K^{3/4+\varepsilon}\right).$
@@
$4\mathop{{\sum}^{\raisebox{-3.0pt}{\makebox[0.0pt][l]{$\flat$}}}}_{d}\,\phi\left(\frac{|d|}{X}\right)\sum_{k\in\mathbb{Z}}h\left(\frac{2k-1}{K}\right)\sum_{f\in B_{k}}\sum_{m_{1}=1}^{\infty}\sum_{m_{2}=1}^{\infty}\frac{\lambda_{f}(m_{1})\lambda_{f}(m_{2})\chi_{d}(m_{1}m_{2})}{\sqrt{m_{1}m_{2}}}V_{k}\left(\frac{m_{1}}{|d|}\right)V_{k}\left(\frac{m_{2}}{|d|}\right)\chi_{d}(n)\sum_{u|n}c_{n}(u)\lambda_{f}(u).$
@@
$4\mathop{{\sum}^{\raisebox{-3.0pt}{\makebox[0.0pt][l]{$\flat$}}}}_{d}\,\phi\left(\frac{|d|}{X}\right)\sum_{k\in\mathbb{Z}}h\left(\frac{2k-1}{K}\right)\sum_{m_{1}=1}^{\infty}\sum_{m_{2}=1}^{\infty}\frac{\chi_{d}(m_{1}m_{2}n)}{\sqrt{m_{1}m_{2}}}V_{k}\left(\frac{m_{1}}{|d|}\right)V_{k}\left(\frac{m_{2}}{|d|}\right)\sum_{u|n}c_{n}(u)\sum_{j|(m_{1},m_{2})}\sum_{f\in B_{k}}\lambda_{f}\left(\frac{m_{1}m_{2}}{j^{2}}\right)\lambda_{f}(u).$
@@
$4\mathop{{\sum}^{\raisebox{-3.0pt}{\makebox[0.0pt][l]{$\flat$}}}}_{d}\,\phi\left(\frac{|d|}{X}\right)\sum_{k\in\mathbb{Z}}h\left(\frac{2k-1}{K}\right)\sum_{m_{1}=1}^{\infty}\sum_{m_{2}=1}^{\infty}\frac{\chi_{d}(m_{1}m_{2}n)}{\sqrt{m_{1}m_{2}}}V_{k}\left(\frac{m_{1}}{|d|}\right)V_{k}\left(\frac{m_{2}}{|d|}\right)\sum_{u|n}c_{n}(u)\sum_{j|(m_{1},m_{2})}1_{m_{1}m_{2}/j^{2}=u}+\text{error term},$
@@
$4\mathop{{\sum}^{\raisebox{-3.0pt}{\makebox[0.0pt][l]{$\flat$}}}}_{d}\,\phi\left(\frac{|d|}{X}\right)\sum_{k\in\mathbb{Z}}h\left(\frac{2k-1}{K}\right)\sum_{m_{1}=1}^{\infty}\sum_{m_{2}=1}^{\infty}\frac{\chi_{d}(m_{1}m_{2}n)}{\sqrt{m_{1}m_{2}}}V_{k}\left(\frac{m_{1}}{|d|}\right)V_{k}\left(\frac{m_{2}}{|d|}\right)\sum_{u|n}c_{n}(u)\sum_{j|(m_{1},m_{2})}\mathbb{E}\left(X\left(\frac{m_{1}m_{2}}{j^{2}}\right)X(u)\right)$
@@
Next we will analyse the error term. In what follows we temporarily think $d$ as fixed and postpone the summation over $d$ to the last moment. We execute the $k$-sum using (5.2.1) to see that the off-diagonal equals
$-2\sqrt{\pi}K\mathop{{\sum}^{\raisebox{-3.0pt}{\makebox[0.0pt][l]{$\flat$}}}}_{d}\phi\left(\frac{|d|}{X}\right)\sum_{m_{1}=1}^{\infty}\sum_{m_{2}=1}^{\infty}\frac{\chi_{d}(m_{1})\chi_{d}(m_{2})}{m_{1}^{3/4}m_{2}^{3/4}}\sum_{u|n}c_{n}(u)u^{-1/4}$
$\qquad\times\textnormal{Im}\left(e^{-2\pi i/8}\sum_{j|(m_{1},m_{2})}\sum_{c=1}^{\infty}\frac{S(m_{1}m_{2}/j^{2},u;c)}{\sqrt{c}}\cdot e\left(\frac{2\sqrt{m_{1}m_{2}u}}{jc}\right)W_{K}^{(2)}\left(\frac{m_{1}}{|d|},\frac{m_{2}}{|d|},\frac{K^{2}cj}{8\pi\sqrt{m_{1}m_{2}u}}\right)\right)+O\left(X^{3+\varepsilon}/K^{2}\right)$
$=-2\sqrt{\pi}K\mathop{{\sum}^{\raisebox{-3.0pt}{\makebox[0.0pt][l]{$\flat$}}}}_{d}\phi\left(\frac{|d|}{X}\right)\sum_{m_{1}=1}^{\infty}\sum_{m_{2}=1}^{\infty}\frac{\chi_{d}(m_{1})\chi_{d}(m_{2})}{m_{1}^{3/4}m_{2}^{3/4}}\sum_{u|n}c_{n}(u)u^{-1/4}$
(9.1) $\qquad\times\textnormal{Im}\left(e^{-2\pi i/8}\sum_{(j,d)=1}\frac{1}{j}\sum_{c=1}^{\infty}\frac{S(m_{1}m_{2},u;c)}{\sqrt{c}}\cdot e\left(\frac{2\sqrt{m_{1}m_{2}u}}{c}\right)W_{K}^{(2)}\left(\frac{m_{1}j}{|d|},\frac{m_{2}j}{|d|},\frac{K^{2}c}{8\pi\sqrt{m_{1}m_{2}u}}\right)\right)+O\left(X^{3+\varepsilon}/K^{2}\right),$

### v2

@@ 6.1. Random model
We need a probabilistic random model for the Hecke eigenvalues $\lambda_{f}(m)$, which we now introduce following [19]. To motivate the model, let $G^{\sharp}$ denote the set of conjugacy classes of $G=\mathrm{SU}(2)$ endowed with the Sato–Tate measure. Note that if $m>1$ has the prime factorisation $m=p_{1}^{\alpha_{1}}\cdots p_{\ell}^{\alpha_{\ell}}$, then we have
@@
with $\alpha_{p},\beta_{p}$ being the Satake parameters of $f$ at a prime $p$. We write $g_{f}^{\sharp}(p)$ for the conjugacy class of the matrix $g_{f}(p)$. Then it can be shown that for a fixed set of primes $\{p_{1},...,p_{\ell}\}$, the $\ell$-tuple $\{g_{f}^{\sharp}(p_{1}),...,g_{f}^{\sharp}(p_{\ell})\}_{f\in B_{k}}$ of conjugacy classes equidistributes inside $(G^{\sharp})^{\ell}$ as $k\longrightarrow\infty$.
This suggests the following random model for $\lambda_{f}(m)$. Let $\{g_{p}^{\sharp}\}_{p}$ be a sequence of independent random variables with values in $G^{\sharp}$ that are distributed according to the Sato–Tate measure. Then we define $X(1)=1$ and
@@
and $\mathbb{E}(X(m))=1_{m=1}$. By linearity of the expectation these give
(6.2) $\mathbb{E}(X(m)X(n))=\sum_{\begin{subarray}{c}d|(m,n)\\ mn=d^{2}\end{subarray}}1=1_{m=n}.$
@@ 6.2. Construction of a random mollifier
$M_{1}(X;d)=(\log K)^{1/2}\sum_{n\leq K^{2\delta_{0}}}\frac{\lambda(n)\chi_{d}(n)h_{1}(n)}{2^{\Omega(n)}\sqrt{n}}\sum_{u|n}c_{n}(u)X(u).$
@@ 7. Character sum
Let $d$ be an odd squarefree integer and let $c,u,v,$ and $\eta$ be natural numbers so that $v\eta=u[c,d]^{2}/c^{2}$. Then the sum
$\sum_{x\,([c,d])}\sum_{w\,([c,d])}\chi_{d}(x)\chi_{d}(w)S\left(xw,u;c\right)e\left(\frac{xv+w\eta}{[c,d]}\right)$
@@
(7.1) $\mathop{{\sum}^{\makebox[0.0pt][l]{$\star$}}}_{\gamma\,(c)}\sum_{w\,([c,d])}\chi_{d}(w)e\left(\frac{u\overline{\gamma}}{c}+\frac{w\eta}{[c,d]}\right)\sum_{x\,([c,d])}\chi_{d}(x)e\left(\frac{xw\gamma}{c}+\frac{xv}{[c,d]}\right).$
@@
Suppose that $n$ and $u$ are odd positive integers with $nu\leq X^{1-\varepsilon}$. Let $\Phi$ denote a smooth and compactly supported function in $\mathbb{R}^{+}$. Then
$\mathop{{\sum}^{\raisebox{-3.0pt}{\makebox[0.0pt][l]{$\flat$}}}}_{d}\chi_{d}(nu)\Phi\left(\frac{|d|}{X}\right)V_{k}\left(\frac{u}{|d|}\right)=\frac{X\pi^{2}}{9}\left(\int\limits_{0}^{\infty}\Phi(\xi)\,\mathrm{d}\xi\right)\prod_{p|nu}\left(1+\frac{1}{p}\right)^{-1}\cdot 1_{nu=\square}+O_{\varepsilon}\left(X^{3/4}k^{-1/4+\varepsilon}+X^{1/2+\varepsilon}\sqrt{nu}\right).$
@@
Let us set $Y:=X^{1/2-\varepsilon}/\sqrt{nu}$. We split the $\alpha$-sum in (8.2) into two parts corresponding to $\alpha\leq Y$ and $\alpha>Y$. For the latter terms we estimate the $r$-sum trivially and get the upper bound (recall that $nu\leq X^{1-\varepsilon}$ and $r\alpha^{2}\asymp X$ so $V_{k}(u/r\alpha^{2})\ll 1$ uniformly in $k$ in this range)
For the terms with $\alpha\leq Y$ we will evaluate the $r$-sum by applying Lemma 5.7. The terms where $nu$ is a square will contribute the main term in the zero frequency term on the dual side and the rest will give the error term.
@@
(8.3) $\frac{X}{2}\cdot\frac{1}{2\pi i}\int\limits_{(c)}\left(\int\limits_{0}^{\infty}\Phi(\xi)\xi^{s}\,\mathrm{d}\xi\right)\frac{\Gamma(s+k)}{\Gamma(k)}\left(\frac{X}{2\pi}\right)^{s}e^{s^{2}}\sum_{\begin{subarray}{c}\alpha=1\\ (\alpha,2u)=1\end{subarray}}^{\infty}\frac{\mu(\alpha)}{\alpha^{2}}u^{-s}\frac{\varphi(nu)}{nu}\cdot 1_{nu=\square}\,\frac{\mathrm{d}s}{s},$
where we have added back in the terms with $\alpha>Y$ at the cost of an error term of size $\ll_{\varepsilon}X^{1/2+\varepsilon}\sqrt{nu}$.
@@
(8.5) $\frac{X}{2}\prod_{p\neq 2}\left(1-\frac{1}{p^{2}}\right)\prod_{p|nu}\left(1+\frac{1}{p}\right)^{-1}\cdot 1_{nu=\square}\cdot\frac{1}{2\pi i}\int\limits_{(c)}\left(\int\limits_{0}^{\infty}\Phi(\xi)\xi^{s}\,\mathrm{d}\xi\right)\frac{\Gamma(s+k)}{\Gamma(k)}\left(\frac{X}{2\pi}\right)^{s}e^{s^{2}}u^{-s}\,\frac{\mathrm{d}s}{s}.$
@@
$\frac{X\pi^{2}}{9}\left(\int\limits_{0}^{\infty}\Phi(\xi)\,\mathrm{d}\xi\right)\prod_{p|nu}\left(1+\frac{1}{p}\right)^{-1}\cdot 1_{nu=\square}+O_{\varepsilon}\left(X^{3/4}k^{-1/4+\varepsilon}\right),$
@@
It remains to bound the contribution coming from the other terms (i.e. $\ell\neq 0$) after applying Lemma 5.7. Let
@@
The sum we have to estimate after an application of Poisson summation is
@@
$=(\log K)^{1/2}\sum_{k\in\mathbb{Z}}h\left(\frac{2k-1}{K}\right)\sum_{f\in B_{k}}\omega_{f}\mathop{{\sum}^{\raisebox{-3.0pt}{\makebox[0.0pt][l]{$\flat$}}}}_{d}\,L\left(\frac{1}{2},f\otimes\chi_{d}\right)\phi\left(\frac{|d|}{X}\right)\sum_{n\leq K^{2\delta_{0}}}\frac{h_{1}(n)a_{f}(n)\lambda(n)\chi_{d}(n)}{2^{\Omega(n)}\sqrt{n}}.$
Let us first use the approximate functional equation and then execute the $f$-sum. For the latter, note that using the definition of $a_{f}(n)$ and applying the Petersson formula we have
@@
where the error term contributes
@@
$\qquad\qquad\times\sum_{n\leq K^{2\delta_{0}}}\frac{h_{1}(n)\lambda(n)\chi_{d}(n)}{2^{\Omega(n)}\sqrt{n}}\sum_{u|n}c_{n}(u)\sum_{c=1}^{\infty}\frac{S(m,u;c)}{c}J_{2k-1}\left(\frac{4\pi\sqrt{mu}}{c}\right)$
to the original sum. This gives a negligible contribution by estimating trivially using Weil’s bound for Kloosterman sums and the bound (5.16) as $mu\leq mn\ll K^{1+2\theta_{J}+\varepsilon}X$ by recalling that $J$ was chosen so that $\eta_{2}\leq\theta_{J}\leq e\eta_{2}$ for some very small $\eta_{2}>0$ as well as $X\ll\sqrt{K}$.
The main term contributes
$(\log K)^{1/2}\sum_{k\in\mathbb{Z}}h\left(\frac{2k-1}{K}\right)\mathop{{\sum}^{\raisebox{-3.0pt}{\makebox[0.0pt][l]{$\flat$}}}}_{d}\phi\left(\frac{|d|}{X}\right)\sum_{n\leq K^{2\delta_{0}}}\frac{h_{1}(n)\lambda(n)\chi_{d}(n)}{2^{\Omega(n)}\sqrt{n}}\sum_{u|n}c_{n}(u)\frac{\chi_{d}(u)}{\sqrt{u}}V_{k}\left(\frac{u}{|d|}\right)$
to the original sum.
@@
after executing the $k$-sum using Poisson summation as the error term from Lemma 8.1 contributes $\ll_{\varepsilon}K^{3/4+\varepsilon}X^{3/4}+K^{1-\varepsilon}X^{1/2}=o(XK)$.
@@
$\qquad\qquad=\prod_{p\in I_{j}}\left(1-\frac{\iota(p)\nu_{2}(p)}{2\sqrt{p}}\left(1+\frac{1}{p}\right)^{-1}+\frac{\iota(p^{2})\nu_{2}(p^{2})}{4p}\left(1+\frac{1}{p}\right)^{-1}+O\left(\frac{1}{p^{3/2}}\right)\right).$
Combining the previous estimates with the fact that (which follows from $\iota(p)=1/\sqrt{p}$)
$\prod_{p\in I_{j}}\left(1-\frac{\iota(p)\nu_{2}(p)}{2\sqrt{p}}\left(1+\frac{1}{p}\right)^{-1}+\frac{\iota(p^{2})\nu_{2}(p^{2})}{4p}\left(1+\frac{1}{p}\right)^{-1}+O\left(\frac{1}{p^{2}}\right)\right)^{-1}\ll 1_{j=0}\cdot(\log K)^{O(1)}+1$
@@
$XK(\log K)^{1/2}\prod_{c<p\leq K^{\theta_{J}}}\left(1-\frac{\iota(p)\nu_{2}(p)}{2\sqrt{p}}\left(1+\frac{1}{p}\right)^{-1}+\frac{\iota(p^{2})\nu_{2}(p^{2})}{4p}\left(1+\frac{1}{p}\right)^{-1}+O\left(\frac{1}{p^{2}}\right)\right)$
@@
To estimate the latter product from below note that
@@
as $\eta_{2}$ was chosen to be sufficiently small.
@@
$\prod_{c<p<K^{\theta_{J}}}\left(1-\frac{1}{2p}+O\left(\frac{1}{p^{2}}\right)\right)\gg(\log K)^{-1/2}$
as $\iota(p)=1/\sqrt{p}$, $\iota(p^{2})=1+1/p$, and $\nu_{2}(p)=\nu_{2}(p^{2})=2$.
@@ 9. Relating the Fourth moment to a random model
Let $n\leq K^{4\delta_{0}}$ be an odd positive integer. Then we have
@@
$=\mathop{{\sum}^{\raisebox{-3.0pt}{\makebox[0.0pt][l]{$\flat$}}}}_{d}\phi\left(\frac{|d|}{X}\right)\chi_{d}(n)\sum_{k\in\mathbb{Z}}h\left(\frac{2k-1}{K}\right)\mathbb{E}\left(L(X;d,k)\sum_{u|n}c_{n}(u)X(u)\right)$
$\qquad+2K\widehat{h}(0)\mathop{{\sum}^{\raisebox{-3.0pt}{\makebox[0.0pt][l]{$\flat$}}}}_{d}\,\frac{\varphi(d)}{d}\phi\left(\frac{|d|}{X}\right)\sum_{u|n}\frac{c_{n}(u)}{\sqrt{u}}\chi_{d}(nu)\cdot\frac{1}{2\pi i}\int\limits_{(\frac{1}{2}+\varepsilon)}\frac{e^{2x^{2}}}{x^{2}}\cdot u^{-x}\sum_{v|u}v^{2x}\,\mathrm{d}x+O\left(X^{3/4}K^{1+\varepsilon}\right).$
@@
$4\mathop{{\sum}^{\raisebox{-3.0pt}{\makebox[0.0pt][l]{$\flat$}}}}_{d}\,\phi\left(\frac{|d|}{X}\right)\sum_{k\in\mathbb{Z}}h\left(\frac{2k-1}{K}\right)$
$\qquad\qquad\times\sum_{f\in B_{k}}\sum_{m_{1}=1}^{\infty}\sum_{m_{2}=1}^{\infty}\frac{\lambda_{f}(m_{1})\lambda_{f}(m_{2})\chi_{d}(m_{1}m_{2}n)}{\sqrt{m_{1}m_{2}}}V_{k}\left(\frac{m_{1}}{|d|}\right)V_{k}\left(\frac{m_{2}}{|d|}\right)\sum_{u|n}c_{n}(u)\lambda_{f}(u).$
@@
$4\mathop{{\sum}^{\raisebox{-3.0pt}{\makebox[0.0pt][l]{$\flat$}}}}_{d}\,\phi\left(\frac{|d|}{X}\right)\sum_{k\in\mathbb{Z}}h\left(\frac{2k-1}{K}\right)$
$\qquad\qquad\times\sum_{m_{1}=1}^{\infty}\sum_{m_{2}=1}^{\infty}\frac{\chi_{d}(m_{1}m_{2}n)}{\sqrt{m_{1}m_{2}}}V_{k}\left(\frac{m_{1}}{|d|}\right)V_{k}\left(\frac{m_{2}}{|d|}\right)\sum_{u|n}c_{n}(u)\sum_{j|(m_{1},m_{2})}\sum_{f\in B_{k}}\lambda_{f}\left(\frac{m_{1}m_{2}}{j^{2}}\right)\lambda_{f}(u).$
@@
$4\mathop{{\sum}^{\raisebox{-3.0pt}{\makebox[0.0pt][l]{$\flat$}}}}_{d}\,\phi\left(\frac{|d|}{X}\right)\sum_{k\in\mathbb{Z}}h\left(\frac{2k-1}{K}\right)$
$\qquad\qquad\times\sum_{m_{1}=1}^{\infty}\sum_{m_{2}=1}^{\infty}\frac{\chi_{d}(m_{1}m_{2}n)}{\sqrt{m_{1}m_{2}}}V_{k}\left(\frac{m_{1}}{|d|}\right)V_{k}\left(\frac{m_{2}}{|d|}\right)\sum_{u|n}c_{n}(u)\sum_{j|(m_{1},m_{2})}1_{m_{1}m_{2}/j^{2}=u}+\text{error term},$
@@
$4\mathop{{\sum}^{\raisebox{-3.0pt}{\makebox[0.0pt][l]{$\flat$}}}}_{d}\,\phi\left(\frac{|d|}{X}\right)\sum_{k\in\mathbb{Z}}h\left(\frac{2k-1}{K}\right)$
$\qquad\qquad\times\sum_{m_{1}=1}^{\infty}\sum_{m_{2}=1}^{\infty}\frac{\chi_{d}(m_{1}m_{2}n)}{\sqrt{m_{1}m_{2}}}V_{k}\left(\frac{m_{1}}{|d|}\right)V_{k}\left(\frac{m_{2}}{|d|}\right)\sum_{u|n}c_{n}(u)\sum_{j|(m_{1},m_{2})}\mathbb{E}\left(X\left(\frac{m_{1}m_{2}}{j^{2}}\right)X(u)\right)$
@@
Next we will analyse the error term. We start by executing the $k$-sum using (5.2.1) to see that the off-diagonal equals
$-2\sqrt{\pi}K\mathop{{\sum}^{\raisebox{-3.0pt}{\makebox[0.0pt][l]{$\flat$}}}}_{d}\phi\left(\frac{|d|}{X}\right)\chi_{d}(n)\sum_{m_{1}=1}^{\infty}\sum_{m_{2}=1}^{\infty}\frac{\chi_{d}(m_{1})\chi_{d}(m_{2})}{m_{1}^{3/4}m_{2}^{3/4}}\sum_{u|n}c_{n}(u)u^{-1/4}$
$\qquad\times\textnormal{Im}\left(e^{-2\pi i/8}\sum_{j|(m_{1},m_{2})}\sum_{c=1}^{\infty}\frac{S(m_{1}m_{2}/j^{2},u;c)}{\sqrt{c}}\cdot e\left(\frac{2\sqrt{m_{1}m_{2}u}}{jc}\right)W_{K}^{(2)}\left(\frac{m_{1}}{|d|},\frac{m_{2}}{|d|},\frac{K^{2}cj}{8\pi\sqrt{m_{1}m_{2}u}}\right)\right)$
$\qquad\qquad\qquad\qquad+O\left(X^{3+\varepsilon}/K^{2}\right)$
$=-2\sqrt{\pi}K\mathop{{\sum}^{\raisebox{-3.0pt}{\makebox[0.0pt][l]{$\flat$}}}}_{d}\phi\left(\frac{|d|}{X}\right)\chi_{d}(n)\sum_{m_{1}=1}^{\infty}\sum_{m_{2}=1}^{\infty}\frac{\chi_{d}(m_{1})\chi_{d}(m_{2})}{m_{1}^{3/4}m_{2}^{3/4}}\sum_{u|n}c_{n}(u)u^{-1/4}$
(9.1) $\qquad\times\textnormal{Im}\left(e^{-2\pi i/8}\sum_{(j,d)=1}\frac{1}{j}\sum_{c=1}^{\infty}\frac{S(m_{1}m_{2},u;c)}{\sqrt{c}}\cdot e\left(\frac{2\sqrt{m_{1}m_{2}u}}{c}\right)W_{K}^{(2)}\left(\frac{m_{1}j}{|d|},\frac{m_{2}j}{|d|},\frac{K^{2}c}{8\pi\sqrt{m_{1}m_{2}u}}\right)\right)$
$\qquad\qquad\qquad\qquad+O\left(X^{3+\varepsilon}/K^{2}\right),$